# Political Interventions to Reduce Single-Use Plastics (SUPs) and Price Effects: An Event Study for Austria and Germany

### v1

@@ 5 Discussion and Policy Implications
Price signals show up when the charge is clear and Austria’s ARA Online and Germany’s DIVID improve monitoring (BMK,, 2024; UBA,, 2024). The magnitudes align with theory and with other environmental charges where modest per-unit prices moved behavior, such as bag levies and deposit–refunds (Convery et al.,, 2007; Homonoff,, 2018; Walls,, 2011).
Distributional incidence and substitution margins matter. Flat per-unit charges can be regressive if lower-income households consume more of the targeted items; complementary policies (reusable alternatives, information, or targeted vouchers) can ease burdens while preserving incentives (Fullerton and Metcalf,, 2001).
On control groups during COVID-19. Because the pandemic pushed up prices broadly and re-shaped supply chains, proper control groups are vital. Controls that are not direct substitutes for treated SUP items are, in general, better in this setting because direct substitutes can be simultaneously affected by substitution away from treated goods and by shared shocks, risking bias in the contrasts.
Limitations.
We do not observe quantities, so we cannot trace welfare or litter outcomes directly. Future work can pair prices with volumes and municipality-level litter metrics to measure elasticities and environmental returns.
@@ 6 Conclusion
Producer charges aimed at SUPs show up in retail prices in Austria, both at the product level and in pooled samples. The full sample including Germany’s fund shows a slower and less precise signal in the short run, consistent with its design and timing. Our empirical approach uses two-way clustering for inference, with product and retailer fixed effects and 3-month event-time bins anchored at February 2022. A plausible suggestion for future research is to quantify changes in quantities, substitution to reusables and net environmental benefits.
@@
Appendix A The data pipeline, sample, and control selection
A.1 Notation and operators
@@
$\texttt{sup}(name):\text{ TRUE iff }name\ \texttt{ILIKE}\ p\text{ for some pattern }p\in\mathcal{P}.$
Base relations (as abbreviated).
$\begin{array}[]{ll}\textbf{Prod}(\textit{prod\_id},\ name,\ born\_ts,\ \ldots)&\text{product metadata}\\ \textbf{Off}(\textit{offer\_id},\ \textit{prod\_id},\ \textit{ret\_id},\ ts,\ price,\ \ldots)&\text{offers (high-frequency)}\\ \textbf{Clk}(\textit{prod\_id},\ \textit{ret\_id},\ ts,\ clicks,\ \ldots)&\text{click logs (high-frequency)}\\ \textbf{Ret}(\textit{ret\_id},\ ret\_name,\ ts,\ \ldots)&\text{retailer attributes over time}\end{array}$
A.2 Cohort and event window
Let $T^{\ast}{=}1{,}577{,}836{,}800$ (Unix; 2020-01-01). Products born after $T^{\ast}$:
@@ A.2 Cohort and event window
Attach time variables to offers and clicks:
@@
Keep event months in [24,36][-24,36] and cohort products:
$\textbf{Off}^{win}$ $=\sigma_{-24\leq\text{e}\leq 36}(\textbf{Off}^{t})\ \bowtie_{\text{prod\_id}}\ \pi_{\text{prod\_id}}(\textbf{Prod}^{+}),$ (4)
$\textbf{Clk}^{win}$ $=\sigma_{-24\leq\text{e}\leq 36}(\textbf{Clk}^{t})\ \bowtie_{\text{prod\_id}}\ \pi_{\text{prod\_id}}(\textbf{Prod}^{+}).$
A.3 Stable retailer snapshot (last observation)
Let $t^{\star}(\text{ret\_id})$ be the last timestamp in Ret for a retailer:
@@ A.3 Stable retailer snapshot (last observation)
A.4 Treatment and controls
Define the SUP predicate via the pattern set $\mathcal{P}$ in Tbl. 3.
@@ A.4 Treatment and controls
Table 3: SUP keyword classes (examples of ILIKE patterns).
@@
Beverage cups %to-go becher%, %einwegbecher%, %kunststoffbecher%,
%getränkebecher%
@@
A.5 Monthly aggregation and index
Aggregate to $(i{=}\text{prod\_id},j{=}\text{ret\_id},t{=}\text{month})$:
@@ A.5 Monthly aggregation and index
Attach event bins ($b{:=}\lfloor e/3\rfloor$):
@@
Compute base prices at $m_{0}$ and form an index:
@@
A.6 Data analysis sample splits
Join clicks and retailer snapshot:
@@ A.6 Data analysis sample splits
This is the analysis table (plastics_regulation_obs).
Define treated and control panels passed to estimation:
@@
Export the variables used in estimation:
@@
A.7 Estimation procedure
For each sample (treated or control), the regression estimated in R is
$Y_{ijt}\sim\sum_{b\neq 0}\mathbf{1}\{b_{t}=b\}\;\Big|\;\text{FE}_{i}+\text{FE}_{j},\quad\text{clusters: }(i,j),$ (13)
corresponding to fixest::feols( Y ~ i(event_bin, ref=0) | prod_id+ret_name, cluster ~ prod_id+ret_name). We report figures for $Y=P$ (level index) with $90\%$ CIs and windowed DiD contrasts (6M, 12M, full post–pre) as linear combinations of bin coefficients.
A.8 Timing and interpretation
Producer charges in Austria begin in 2022; Germany’s fund begins in 2024. Because $P=100\cdot\overline{p}/\overline{p}_{0}$ with base $m_{0}{=}$February 2022, bin coefficients $\beta_{b}$ are percentage-point deviations from that base month.

### v2

@@ 5 Discussion and Policy Implications
5.1 Price incidence and control groups
Clear price responses appear in Austrian retail data when producer charges are precisely defined and supported by established reporting systems (e.g. ARA Online in Austria and DIVID in Germany) (BMK,, 2024; UBA,, 2024). The estimated magnitudes for pooled SUP products and for balloons are consistent with prior evidence on taxes and charges applied to energy products and plastic bags (Marion and Muehlegger,, 2011; Besley and Rosen,, 1999; Convery et al.,, 2007; Homonoff,, 2018). In these earlier settings, relatively small per-unit charges translated into visible retail price adjustments and meaningful changes in economic incentives.
Control group construction plays a central role in this analysis. The primary control set consists of non-SUP products that share retail channels, digital platform exposure, and broad consumption environments with treated items. A separate strict control group based on products such as graphics cards contrasts SUP items with goods subject to very different upstream cost dynamics, including semiconductor supply cycles and cryptocurrency-driven demand. Estimates based on this strict set are best interpreted as a stress test for sector-specific pressures rather than as the main counterfactual path for SUP packaging. The substantive findings therefore rely on the broader non-SUP controls and on within-country comparisons.
The sample period coincides with substantial macroeconomic disturbances, including the COVID-19 pandemic and the war in Ukraine. Energy and feedstock prices rose sharply from 2022 onward, affecting plastics manufacturing costs across Europe (Packaging Europe,, 2022; Plastics Europe,, 2022). The empirical design addresses these forces through three features: (i) normalization of prices to a common base month; (ii) use of non-SUP controls that share retailers and platform-level exposure; and (iii) examination of pre-event bins for similar price movements. These steps do not rule out all sources of confounding, and part of the estimated price response may reflect differences in how SUP and non-SUP sectors react to shared cost shocks. This possibility informs the interpretation of the results.
5.2 Germany, the war in Ukraine, and macroeconomic context
For Germany, the post-policy window remains limited. The Single-Use Plastics Fund took effect in 2024, and the sample ends in May 2025, leaving only a small number of three-month event bins fully within the post-charge period. Over the same interval, German retail prices responded to war-related energy shocks and ongoing supply-chain adjustments affecting plastics and packaging costs (Packaging Europe,, 2022; Plastics Europe,, 2022). These concurrent forces complicate attempts to isolate the price effects of the fund in cross-country comparisons.
The German estimates are therefore presented as preliminary and descriptive. They provide an early view of price movements in treated segments relative to controls after the fund’s introduction but do not support precise quantitative statements about pass-through. A more complete assessment would require longer post-policy panels, broader coverage of German SUP products, and possibly additional macroeconomic series. Carrying out such an analysis would involve further sampling of the underlying database and the incorporation of external data sources, which falls outside the scope of the present study.
5.3 Quantities, welfare, and future research
The database used in this study contains detailed information on retail prices but does not observe quantities or sales volumes. The analysis therefore addresses only the first step in the policy transmission chain: whether producer charges and fund payments associated with single-use plastics (SUPs) are passed through to shelf prices. It does not address the subsequent steps—whether higher prices reduce SUP consumption, induce substitution toward reusable alternatives, or ultimately lower litter and environmental damages.
A simple CES model and testable predictions.
To clarify the key margins, consider a representative consumer who allocates expenditure between a SUP composite good $q_{S}$ and a reusable (or otherwise non-SUP) composite $q_{R}$ with CES preferences,
$U(q_{S},q_{R})=\left[\omega q_{S}^{\frac{\sigma-1}{\sigma}}+(1-\omega)q_{R}^{\frac{\sigma-1}{\sigma}}\right]^{\frac{\sigma}{\sigma-1}},\qquad\sigma>0,$ (2)
subject to the budget constraint $P_{S}q_{S}+P_{R}q_{R}=I$, where $P_{S}$ and $P_{R}$ are price indices and $I$ is income. The elasticity of substitution between SUP and non-SUP composites is $\sigma$. Standard CES demand implies
$q_{S}=\omega\left(\frac{P_{S}}{P}\right)^{-\sigma}\frac{I}{P_{S}},\qquad q_{R}=(1-\omega)\left(\frac{P_{R}}{P}\right)^{-\sigma}\frac{I}{P_{R}},$ (3)
where $P=\left[\omega P_{S}^{1-\sigma}+(1-\omega)P_{R}^{1-\sigma}\right]^{\frac{1}{1-\sigma}}$ is the CES price index.
On the supply side, let the pre-policy marginal cost of a representative SUP product be $c_{S}$, and suppose a per-unit producer charge $\tau$ increases marginal cost to $c_{S}+\tau$. Let $\rho\in[0,1]$ denote the pass-through rate into the consumer price (under competition, market power, and/or nominal rigidities), so that
$\Delta P_{S}=\rho\,\tau.$ (4)
The empirical analysis in this paper speaks directly to (4) by estimating changes in $P_{S}$ for SUP products relative to non-SUP products.
This simple structure yields three predictions that can be taken to data once quantities (or adequate proxies) are available:
  1. 1.
    Pass-through: a positive producer charge raises retail SUP prices, $\Delta P_{S}>0$, with magnitude governed by $\rho$.
  2. 2.
    Demand response and substitution: higher SUP prices reduce SUP demand according to the CES elasticity, with (semi-)elasticity
    $\frac{\partial\ln q_{S}}{\partial\ln P_{S}}=-\sigma-(1-\text{s}_{S})\sigma,\quad\text{and approximately}\quad\Delta\ln q_{S}\approx-\sigma\,\Delta\ln\!\left(\frac{P_{S}}{P_{R}}\right)\ \ \text{for small changes},$ (5)
    where $\text{s}_{S}$ is the expenditure share on $q_{S}$. [^1] The exact own-price elasticity under CES depends on how the aggregate price index $P$ moves with $P_{S}$. For empirically relevant small price changes, the approximation in relative prices is often a convenient sufficient statistic.
  3. 3.
    Environmental impact: if environmental damages are proportional to SUP quantity, $D=\delta q_{S}$ with $\delta>0$, then the policy reduces damages whenever it reduces $q_{S}$, with $\Delta D=\delta\,\Delta q_{S}<0$ if $\Delta q_{S}<0$.
Welfare accounting.
To connect the model to welfare, let total welfare be
$W=CS+\Pi+T-D,$ (6)
where $CS$ denotes consumer surplus, $\Pi$ producer profits, $T$ net public revenue from the charge (or fund payments), and $D$ environmental damages. A per-unit charge affects welfare through: (i) incidence (how much of $\tau$ is borne by consumers vs. producers, governed by $\rho$), (ii) the quantity response (governed by $\sigma$), and (iii) the conversion of reduced SUP use into environmental benefits (governed by $\delta$ and the mapping from SUP consumption to litter or waste).
In the CES setting, a convenient sufficient-statistics representation for the efficiency component of welfare change from a small tax increase is the Harberger triangle relative to the external benefit. For a marginal change $d\tau$,
$dW\approx\underbrace{\delta\,(-dq_{S})}_{\text{environmental benefit}}\;-\;\underbrace{\tfrac{1}{2}\,dq_{S}\,dP_{S}}_{\text{DWL from reduced consumption}},\qquad dP_{S}=\rho\,d\tau,$ (7)
where the demand slope (hence $dq_{S}$) is pinned down by $\sigma$ and baseline shares via (3)–(5). This expression highlights why price pass-through alone is not sufficient for welfare analysis: without $q_{S}$ (or credible proxies), one cannot quantify either the deadweight loss term or the environmental benefit term.
Implications for the present evidence.
The estimates in this paper provide evidence of positive pass-through ($\rho>0$) in Austria for several SUP segments, consistent with Prediction 1. However, Predictions 2–3 and the welfare decomposition in (7) require data on quantities and substitution patterns. In particular, even large price effects may translate into modest environmental benefits if $\sigma$ is small, whereas smaller price effects could yield substantial benefits if substitution away from SUPs is highly elastic.
Future research direction using modern DiD/event-study designs.
A natural next step is to construct linked panels that add (i) product-level sales volumes, or (ii) credible proxies for quantities (e.g. scanner data, retailer transaction aggregates, or platform-level order counts), and/or (iii) municipal outcomes such as litter counts, waste collection volumes, or recycling rates. With such data, the research designs emphasized in Bertrand et al., (2004); Callaway and Sant’Anna, (2021); Sun and Abraham, (2021) can be deployed to strengthen causal interpretation in three ways.
[…]
Implementing these extensions would require new sampling across the large product taxonomy, careful keyword design, and harmonized product classifications across countries. The present study prepares the ground by documenting a transparent pipeline and identification strategy and by clarifying where macroeconomic shocks and the absence of quantities currently limit welfare and environmental conclusions.
@@ 6 Conclusion
This paper studies price responses to producer charges targeting single-use plastics using high-frequency retail data from Austria and Germany. For Austria, the results show price pass-through both at the individual product level and in pooled analyses. Prices for balloons and for the pooled SUP group rise relative to non-SUP controls when measured as index values with February 2022 as the base month. For plastic to-go cups, price movements in the first year after the charge are more volatile, and while the full-period contrast is positive, it is estimated with limited precision.
For Germany, the evidence should be read as early and descriptive. The Single-Use Plastics Fund took effect only in 2024, and the sample ends in mid-2025, leaving a short post-policy window. This restricts the ability to assess longer-run price responses. In the pooled cross-country analysis, estimated effects mainly reflect the Austrian experience combined with a brief and variable German post-treatment period.
The empirical design relies on a fixed-effects event-study specification with three-month event-time bins and two-way clustered standard errors. The approach remains intentionally parsimonious and stays within the standard difference-in-differences framework, reflecting the structure of the available product–retailer panel and the uneven length of price spells. Extensions that allow for richer forms of treatment effect heterogeneity, alternative weighting schemes, or staggered adoption at finer levels would require longer post-policy panels and additional data. Once such data become available, recent advances in the difference-in-differences literature provide natural tools for further analysis (Bertrand et al.,, 2004; Callaway and Sant’Anna,, 2021; Sun and Abraham,, 2021).
@@
Appendix A Data pipeline, sample construction, and control selection
This appendix documents the construction of the analysis dataset, the definition of treatment and control groups, and the aggregation and indexing procedures used in estimation.
A.1 Notation and relational operators
@@
$\texttt{sup}(name):\text{ TRUE if }name\ \texttt{ILIKE}\ p\text{ for some pattern }p\in\mathcal{P}.$
Base relations.
[…]
A.2 Cohort definition and event window
Let $T^{\ast}=1{,}577{,}836{,}800$ denote January 1, 2020 in Unix time. We restrict attention to products entering the database after this date:
@@ A.2 Cohort definition and event window
Time variables are attached to offers and clicks as follows:
@@
We retain observations within the event window [24,36][-24,36] months and restrict to cohort products:
$\textbf{Off}^{win}$ $=\sigma_{-24\leq e\leq 36}(\textbf{Off}^{t})\;\bowtie_{\text{prod\_id}}\;\pi_{\text{prod\_id}}(\textbf{Prod}^{+}),$ (10)
$\textbf{Clk}^{win}$ $=\sigma_{-24\leq e\leq 36}(\textbf{Clk}^{t})\;\bowtie_{\text{prod\_id}}\;\pi_{\text{prod\_id}}(\textbf{Prod}^{+}).$
A.3 Retailer snapshot
Retailer attributes are fixed at their most recent observed values. Let $t^{\star}(\text{ret\_id})$ denote the last timestamp for a retailer:
@@ A.3 Retailer snapshot
A.4 Treatment and control definitions
Products subject to SUP regulation are identified via keyword patterns $\mathcal{P}$ listed in Table 4:
@@ A.4 Treatment and control definitions
Table 4: SUP keyword classes (illustrative ILIKE patterns).
@@
Beverage cups %to-go becher%, %einwegbecher%, %kunststoffbecher%, %getränkebecher%
@@
A.5 Monthly aggregation and price index
Offers and clicks are aggregated to the product–retailer–month level:
@@ A.5 Monthly aggregation and price index
Event time and three-month bins $b=\lfloor e/3\rfloor$ are assigned:
@@
Base prices at $m_{0}$ are used to construct a price index:
@@
A.6 Analysis samples
Clicks and retailer attributes are merged to form the analysis table:
@@ A.6 Analysis samples
Treated and control samples are defined as:
@@
The variables exported to estimation are:
@@
A.7 Estimation and interpretation
For each sample, the estimated regression is:
$Y_{ijt}\sim\sum_{b\neq 0}\mathbf{1}\{b_{t}=b\}\;\Big|\;\text{FE}_{i}+\text{FE}_{j},\qquad\text{clusters: }(i,j),$ (19)
implemented in R via fixest::feols(Y ~ i(event_bin, ref=0) | prod_id + ret_name, cluster ~ prod_id + ret_name). Figures report 90% confidence intervals for $Y=P$ (price index levels) and windowed DiD contrasts (6 months, 12 months, and full post–pre).
A.8 Timing and policy interpretation
Producer charges apply in Austria from 2022 onward, while Germany’s Single-Use Plastics Fund takes effect in 2024. Since $P=100\cdot\overline{p}/\overline{p}_{0}$ with base month February 2022, bin coefficients represent percentage-point deviations from that baseline. DiD contrasts summarize differences between average post- and pre-policy bins in these index points.